# Analog model of Kerr black holes in scalar-tensor theories using binary Bose-Einstein condensates: Quasibound states and tachyonic instabilities

### v1

@@ Abstract
We propose an analog model of the Kerr black hole in scalar-tensor theories using 2-component Bose-Einstein Condensates. We focus on two types of gapped excitations induced by the modes of two-component condensates with a relative phase of 00 and $\pi$, analogous to the massive scalar field with positive and negative mass squared, respectively. The inclusion of space-dependent Rabi coupling can induce an effective space-dependent mass term. We study superradiant instabilities resulting from the quasibound states corresponding to positive mass squared and the tachyonic instabilities arising from negative mass squared in both frequency and time domains. These instabilities resemble two possible mechanisms that make Kerr black holes unstable in scalar-tensor gravity with the presence of matter around the black hole. Our proposed analog gravity model could potentially be implemented in future experiments, drawing from the success of recent analog rotating black hole implementations.
@@ I INTRODUCTION
Analog gravity can be traced back to the pioneering work by Unruh [1]. The idea is to use the sound waves in a moving fluid as an analog to light waves in curved spacetime and further show that supersonic fluid can generate a “dumb hole”, an acoustic analog of a black hole. Since its development, the analog gravity program has received much attention for exploring fundamental physics through interdisciplinary research among particle/astro physicists and condensed matter physicists (for a review see [2]).
@@
The linear perturbations around a perturbed black hole can be probed through its damped oscillatory behavior in the form of so-called quasinormal modes (QNMs) [3, 4, 5]. The QNMs have a complex frequency where its real part gives the frequency of oscillations and the imaginary part determines the lifetime. During the ringdown phase of a perturbed black hole, possibly formed from a binary black hole merger, the observed spectrum of QNMs encodes information such as the angular momentum and mass of a black hole. Importantly, this encoding is independent of the initial conditions of perturbations, which allows for testing both general relativity and alternative theories of relativity [6]. In particular, recent studies have focused on mimicking the geometry of a spinning black hole in 1+2 dimensions using a draining bathtub vortex flow [7, 8, 9, 10]. An analog model of rotating black holes has also been proposed in the photon fluid [11, 12, 13]. With the state-of-the-art experimental technology of Bose-Einstein condensates (BECs), the first successful experiment of Hawking radiation and its time dependence in a single-component BEC system in 1+1 dimensions is implemented in [14, 15]. Subsequently, a theoretical extension to binary BECs in 1+1 dimensions has been put forward as an analog model to simulate the behavior of the massive scalar field in the early universe and black holes [16, 17, 18, 19, 20, 21]. A further extension to 1+2 dimensions provides a fantastic platform to study quantum vortex instabilities in the background of analogous rotating black holes [22, 23].
@@
The most studied alternatives to general relativity are scalar-tensor theories [24]. Recent studies in [25, 26, 27] have revealed the mechanism that renders Kerr black holes unstable due to the presence of matter in the vicinity of the black holes, which alters the QNMs spectrum. One of the mechanisms involves positive mass squared due to the existence of so-called quasibound states, leading to superradiant instabilities when the spin of a black hole exceeds certain threshold. The other comes from negative mass squared, causing tachyonic instabilities that may further trigger the development of black-hole hairs, known as ”spontaneous scalarization”.
@@
In this Letter, we propose an analog model of Kerr black holes in scalar-tensor theories using binary BEC systems. Introducing Rabi transition between two atomic hyperfine states, the system can be represented by a coupled two-field model that involves both gapless excitations and gapped excitations [16, 17, 18, 19, 20]. The advent of experimental studies on the tunable binary BECs in [28, 29, 30, 31] makes it possible to observe these two types of excitations. We primarily focus on the gapped modes, which are analogous to the massive scalar field. In the background of zero and $\pi$ phase difference between two-component condensates, the induced mass term can be positive and negative, respectively [32, 33, 34]. Subsequently, the space-dependent Rabi coherent coupling [31, 35, 36, 37] can generate the spatially dependent mass term, allowing for the simulation of superradiant and tachyonic instabilities in the background of a draining vortex flow with an analogous geometry of rotating black holes.
@@ II FORMALISM
We consider tunable binary BECs with identical atoms in two distinct internal hyperfine states. Experimentally, one can tune the values of scattering length using Feshbach resonances, such as those associated with the two hyperfine states of ${}^{87}\text{Rb}$ [38, 39, 40, 41]. Additionally, we introduce the Rabi transition between these states with the strength determined by the Rabi-frequency $\Omega$, where $\Omega\geq 0$. With the unit $\hbar=1$ throughout this paper, the coupled time-dependent equations of motion in $1+2$ dimensions are expressed by
@@
with $i\neq j$ and $i,j=1,2$. $m_{a}$ represents atomic mass and $V_{i}$ denotes the external potentials on the hyperfine states $i$. Furthermore, $g_{ii},g_{ij}$ give the interaction strengths of atoms between the same hyperfine states and different hyperfine states, respectively. The coupling strengths are related to the scattering lengths. The condensate wavefunctions are given by the expectation value of the field operator $\langle\hat{\Psi}_{i}\rangle$  , $\langle\hat{\Psi}_{i}\rangle(\mathbf{r},t)=\sqrt{\rho_{i}(\mathbf{r},t)}\,e^{i\theta_{i}(\mathbf{r},t)-i\mu t}$ with the chemical potential $\mu$. The condensate flow velocities are given by $\mathbf{v}_{i}={\gradient}\theta_{i}/m_{a}$. The equations for $\rho_{i}$ and $\theta_{i}$ of the condensate wavefunctions can be found in [18, 17, 19]. The perturbations around the stationary wavefunctions are defined through $\hat{\Psi}_{i}=\langle\hat{\Psi}_{i}\rangle(1+\hat{\phi}_{i})\,,$ where the fluctuation fields decompose in terms of the density and the phase as $\hat{\phi}_{i}=\delta{\hat{n}_{i}}+i\delta\hat{\theta}_{i}=\frac{\delta\hat{\rho}_{i}}{2\rho_{i}}+i\delta\hat{\theta}_{i}\,.$ According to [18], one can decouple the equations for the general spatially dependent condensate wavefunctions and the coupling strengths by choosing the time-independent background solutions $\rho_{1}=\rho_{2}=\rho$, $\theta_{1}=\theta_{2}=\theta$ or $\theta_{1}=\theta,\theta_{2}=\theta\pm\pi$, and $g_{11}=g_{22}=g$ (See Supplements) [32, 34]. The chosen scattering parameters of $g_{12}<g$ in the binary systems can result in a miscible state of the background condensates [28, 29, 30, 31].
The decoupled equations are shown in [18, 17, 19] in terms of the (gapless) density and (gapped) spin modes given by $\sqrt{2}\delta\hat{n}_{d/s}=\delta\hat{n}_{1}\pm\delta\hat{n}_{2}$ and $\sqrt{2}\delta\hat{\theta}_{d/s}=\delta\hat{\theta}_{1}\pm\delta\hat{\theta}_{2}$. Combining them yields a single equation for $\delta\hat{\theta}_{s}$ in the form of the Klein-Gordon equation for a massive scalar field,
@@
where the spatial dependent sound speed is given by $c^{2}=[(g-g_{12})\rho\pm\Omega]/m_{a}$, and the effective mass by $m^{2}_{\text{eff}}=\pm 2m_{a}^{2}c\,\Omega/\rho$, with the "+""""+"" + " sign corresponding to the zero-phase difference and the """""-"" - " sign to the $\pi$-phase difference. The acoustic metric $\mathbf{g}^{\mu\nu}$ depends on the background $\rho$ and $\theta$. Following [23], to mimic the geometry of a rotating black hole, we consider the draining vortex with the velocity flow
@@
where the winding number $\ell$ is taken as an integer (hereafter $\ell=1$) for quantum vortices and $d>0$ for the draining flow. In the Thomas-Fermi approximations, the region of interest is far from the core of the vortex with the radius $r\gg R_{\text{TF}}=\sqrt{\frac{d^{2}+l^{2}}{2m_{a}\rho_{\infty}(g+g_{12})}}$, where the density $\rho$ can be treated as a constant, $\rho\simeq\rho_{\infty}$ given by $\rho_{\infty}=\frac{1}{g+g_{12}}\left(\mu-V_{0}\pm\frac{\Omega}{2}\right)$ for an uniform external potential $V_{0}$. The details can be seen in Supplements. Using the coordinate transformations $d\tilde{t}=dt-(c^{2}/v_{r}^{2}-1)^{-1}dr,$ $d\tilde{\phi}=d\phi-\frac{v_{r}v_{\phi}}{r(c^{2}-v_{r}^{2})}dr,$ the metric becomes
@@
$\propto-\left(1-\frac{r_{E}^{2}}{r^{2}}\right)d\tilde{t}^{2}+\left(1-\frac{r^{2}_{H}}{r^{2}}\right)^{-1}dr^{2}-\frac{2\ell}{m_{a}}\,d\tilde{\phi}d\tilde{t}+r^{2}d\tilde{\phi}^{2}\,$ (4)
with $\tilde{x}=(\tilde{t},r,\tilde{\phi})$. Apparently, the acoustic horizon and the ergosphere are located at $r_{\text{H}}=d/m_{a}c$ and $r_{\text{E}}=\sqrt{d^{2}+\ell^{2}}/m_{a}c$, respectively. Hereafter, we simplify the formula by defining dimensionless variables as $r/r_{H}\rightarrow r$, $c\,\tilde{t}/r_{H}\rightarrow{t}$ and $r_{H}\Omega/c\rightarrow\Omega$. Let us now provide the order-of-magnitude of the relevant parameters from the experiment. In the case of Rubidium atoms with a mass $m_{a}\sim 10^{-25}\text{kg}$ and a scattering length $a\sim 6\text{nm}$, the density $\rho_{\infty}$ can be prepared as much as $\sim 10^{25}\text{m}^{-3}$, giving the healing length and $R_{\text{TF}}$ such as $\xi\sim R_{\text{TH}}\sim 10^{2}\text{nm}$. Considering $r_{\text{H}}\sim 10\xi$, from $v_{r}=c$, the value of $d$ is about $d\sim 1$. Also, the measuring time to trace the evolution of the perturbed fields will be limited by the lifetime of the condensates, say $\sim 1\text{s}$, giving the dimensionless time of order ${t}\sim 10^{3}$. Using the phase fluctuation field of the form $\delta\theta_{s}({t},r,\tilde{\phi})=\frac{H_{m}(r,{t})}{\sqrt{r}}\exp\left(im\tilde{\phi}\right)$ in the Klein-Gordon equation, the radial equation in the tortoise coordinate $r^{\ast}=r+\frac{1}{2}\log{\Big{|}\frac{r-1}{r+1}\Big{|}}$ ) reads
@@
×$\times\left[\frac{1}{r^{2}}\left(m^{2}-\frac{1}{4}\right)+\frac{5}{4r^{4}}+\mu_{s}^{2}(r)\right]\Bigg{\}}H_{m}(r,t)=0$ (5)
with $\mu^{2}_{s}(r)=\pm 2d\Omega(r)$. Further writing $H_{m}(r,t)=H_{m}(r,\omega)e^{-i\omega{t}}$ turns the time-dependent equation into the time-independent Schrodinger-like equation $\left[\partial_{r^{\ast}}^{2}+V_{\text{eff}}(r)\right]H_{m}(r,\omega)=0$ where $V_{\text{eff}}(r)=\left({\omega-\frac{m\ell}{r^{2}}}\right)^{2}-\left(\frac{r^{2}-1}{r^{2}}\right)\left[\frac{5}{4r^{4}}+\frac{1}{r^{2}}\left(m^{2}-\frac{1}{4}\right)+\mu^{2}_{s}(r)\right]$. To establish the analog gravity model of the scalar-tensor theory, the space-dependent mass term is introduced [25, 26]. The mass term due to the coupling to the surrounding matter of the black hole can be parametrized in terms of three parameters. One of them is the magnitude of the mass reflecting the coupling strength between the scalar field and matter in the scalar-tensor gravity theories, which can be controlled by the Rabi coupling $\Omega$. The other two are related to the parametrization of the mass distribution. They are the typical distance of the mass distribution $r_{0}$ from the black hole’s horizon and the distribution width $1/\alpha$. The parametrization of the distance $r_{0}$ can be attributed to the radius of the so-called innermost stable circular orbit of the particle moving around a massive object such as a black hole. The radius of the innermost circular motion depends on the mass and angular momentum (spin) of the massive object. The width of the mass distribution $1/\alpha$ is also an important parameter characterizing the surrounding matter of the black holes.
@@
Figure 2: In the case of positive mass squared, the panel (a) shows the frequency for the mode $n=0$ (red) and $n=1$ (blue) as a function of $\Omega_{0}$ with fixed $r_{0}=15$, and the panel (b) shows the corresponding frequency as a function of $r_{0}$ with $\Omega_{0}=0.2$. (c) depicts the effective potential and the normalized radial profile for the first two modes, and (d) presents a series of peaks at the resonant frequencies in the reflection coefficient spectrum. (d) reveals the growing waveform in time domain and the echo time duration. (c)-(e) share the same values of $\Omega_{0}$ and $r_{0}$ marked in the green line in (a) and (b). Note that $m=1$, and $\alpha=1/2$ are set in all figures.
Given the boundary conditions of the pure ingoing wave $e^{-i(\omega-m\ell)r^{\ast}}$ at the horizon and the pure outgoing wave $e^{i\omega r^{\ast}}$ at infinity, the eigenfrequency of the time-independent Schrodinger-like equation can be found with the complex values $\omega=\omega_{R}+i\omega_{I}$. However, there are two types of instabilities of interest with very different dynamics on the timescale $\tau\propto 1/|\omega_{I}|$: superradiant instabilities are driven by an unstable but long-lived mode where $\omega_{R}\gg\omega_{I}$ , while tachyonic instabilities are featured with an almost pure imaginary frequency $\omega_{I}\gg\omega_{R}$ leading to a short-lived mode. In two-component BECs, the effective mass squared term can be parameterized as
@@
mimicking a mass shell surrounding the black hole in the scalar-tensor gravity theory. We also present the case with the uniform mass square term for comparison in Supplements where we use the continued fraction method [42]. In the case of the spatially dependent mass term (6), we calculate the eigenfrequency numerically using the shooting method [43], where the solutions are obtained by integrating the equation from the horizon onwards and from the far distance $r$ backward with appropriate boundary conditions where they continuously vary over $r$. We also find the eigenfrequency with the generalized Pöschl-Teller (PT) semi-analytical methods [44, 45, 46]. The analytically approximated formulas for the eigenfrequency are given below. The instabilities are shown in the time domain by solving the time-dependent equation (5). For technical details, refer to Supplements.
Quasinormal modes and superradiant instabilities. In the case of the quasinormal modes, their spectrum is strongly modified by the presence of the spatially dependent mass shell, which introduces an additional barrier to the rotational barrier given by the azimuthal angular momentum $m$. The frequencies of quasinormal modes for the real and imaginary parts are found in Fig. 1, which are smaller than their counterparts without the mass shell [47, 48]. The corresponding time domain signals are obtained by solving the time-dependent equation (5) for an initial Gaussian function with reference to [49]. According to [49], small perturbations can lead to a relatively large shift of the eigenfrequency from the unperturbed one, but not to a significant change in the time domain (See an example in Supplements). Here, we present the evolution of the radial profiles, which show a significant change due to the nonzero Rabi coupling strength. The echo time (estimated from the shaded time domain around $\sim 2r_{0}$), which is the time for the radial profile to travel towards the potential barrier at $r=r_{0}$ and return to the starting position and the ringdown of the profile at late times are displayed in Fig 1. More interestingly, when $\Omega_{0}$ or $r_{0}$ reaches its respective critical value, the positive mass shell could further induce superradiant instabilities, where a local bound state is shown for the potential ($-V_{\text{eff}}$) in Fig. 2. This is similar to superradiant instabilities induced by the accretion disk around Kerr black hole in scalar-tensor gravity theories [25, 26, 27]. Let us estimate the resonance spectrum via a discrete quantization condition in the Wentzel-Kramers-Brillouin (WKB) analysis given by

### v2

# Acoustic quasibound states and tachyonic instabilities from binary Bose-Einstein condensates

@@ Abstract
We consider two-component Bose-Einstein condensates (BECs) and introduce the BEC vortex in $1+2$ dimensions. We focus on two types of gapped excitations induced by the modes of two-component BECs with relative phases of 00 and $\pi$, analogous to the massive scalar field with positive and negative mass squared, respectively. The inclusion of space-dependent Rabi coupling can induce an effective space-dependent mass term. We study superradiant instabilities resulting from the quasibound states corresponding to positive mass squared and the tachyonic instabilities arising from negative mass squared in both the frequency and time domains. These instabilities resemble two possible mechanisms that make Kerr black holes unstable in scalar-tensor gravity with the presence of matter around the black hole. Our proposed phenomena could potentially be implemented in future experiments, drawing from the success of recent analog rotating black hole implementations.
@@ I INTRODUCTION
The linear perturbations around a perturbed black hole can be probed through its damped oscillatory behavior in the form of so-called quasinormal modes (QNMs) Berti et al. (2009); Dolan et al. (2012a); Konoplya and Zhidenko (2011). The QNMs have a complex frequency, where the real part gives the frequency of oscillations and the imaginary part determines the lifetime. During the ringdown phase of a perturbed black hole, possibly formed from a binary black hole merger, the observed spectrum of QNMs encodes information such as the angular momentum and mass of a black hole. Importantly, this encoding is independent of the initial conditions of perturbations, which allows for testing both general relativity and alternative theories of relativity Yunes et al. (2016). In particular, recent studies have focused on mimicking the geometry of a spinning black hole in 1+2 dimensions using a draining bathtub vortex flow Visser (1998); Richartz et al. (2015); Torres et al. (2017, 2020). Subsequently, a theoretical extension to binary Bose Einstein condensates (BECs) in 1+1 dimensions has been put forward as an analog model to simulate the behavior of the massive scalar field in the early Universe and black holes Fischer and Schützhold (2004); Liberati et al. (2006); Visser and Weinfurtner (2005); Syu et al. (2022); Syu and Lee (2023); Syu et al. (2019). A further extension to 1+2 dimensions provides a fantastic platform to study quantum vortex instabilities in the background of analogous rotating black holes Berti et al. (2023); Patrick et al. (2022).
@@
Black holes are typically not isolated but are at the centers of galaxies or surrounded by matter such as accretion disks and dark matter. Therefore, the spectra that we observe from gravitational waves should be modified by interactions with their surroundings Barausse et al. (2014); Cardoso, Vitor and Maselli, Andrea (2020); McGee et al. (2020). Recent studies in Cardoso et al. (2013a, b); Lingetti et al. (2022) have revealed mechanisms that render Kerr black holes unstable due to the presence of matter in the vicinity of the black holes within the context of scalar-tensor theories, which alters the QNMs spectrum. One of the mechanisms involves positive mass squared due to the existence of so-called quasibound states, leading to superradiant instabilities when the spin of a black hole exceeds a certain threshold. The other comes from negative mass squared, causing tachyonic instabilities that may further trigger the development of black-hole hairs, known as “spontaneous scalarization”.
@@
In this article, we consider binary BEC systems and introduce the BEC vortex. With Rabi transition between two atomic hyperfine states, the system can be represented by a coupled two-field model that involves both gapless excitations and gapped excitations Fischer and Schützhold (2004); Liberati et al. (2006); Visser and Weinfurtner (2005); Syu et al. (2022); Syu and Lee (2023). The advent of experimental studies on the tunable binary BECs in Kim et al. (2020); Cominotti et al. (2022); Hamner et al. (2011, 2013) makes it possible to observe these two types of excitations. We primarily focus on the gapped modes, which are analogous to the massive scalar field. In the background of zero and $\pi$ phase difference between two-component condensates, the induced mass term can be positive and negative, respectively Zibold et al. (2010); Abbarchi et al. (2013); Tommasini et al. (2003). Subsequently, the space-dependent Rabi coherent coupling Hamner et al. (2013); Mitsunaga et al. (2000); Nicklas et al. (2011); Han et al. (2015) can generate a spatially dependent mass term, enabling the simulation of superradiant and tachyonic instabilities in the background of a draining vortex flow with an analogous geometry to rotating black holes. Our work opens up a new area of research to realize the effects of surrounding matter on black holes in a laboratory experiment using feasible spatially dependent Rabi coupling in binary BEC systems.
@@ II FORMALISM
We consider tunable binary BECs with identical atoms in two distinct internal hyperfine states. Experimentally, one can tune the values of scattering length using Feshbach resonances, such as those associated with the two hyperfine states of ${}^{87}\text{Rb}$ Myatt et al. (1997); Hall et al. (1998); Papp et al. (2008); Tojo et al. (2010). Additionally, we introduce the Rabi transition between these states with the strength determined by the Rabi frequency $\Omega$, where $\Omega\geq 0$. With the unit $\hbar=1$ throughout this paper, the coupled time-dependent equations of motion in $1+2$ dimensions are expressed by
@@
with $i\neq j$ and $i,j=1,2$. $m_{a}$ represents atomic mass and $V_{i}$ denotes the external potentials on the hyperfine states $i$. Furthermore, $g_{ii},g_{ij}$ give the interaction strengths of atoms between the same hyperfine states and different hyperfine states, respectively. The coupling strengths are related to the scattering lengths. The condensate wave functions are given by the expectation value of the field operator $\langle\hat{\Psi}_{i}\rangle$ , $\langle\hat{\Psi}_{i}\rangle(\mathbf{r},t)=\sqrt{\rho_{i}(\mathbf{r},t)}\,e^{i\theta_{i}(\mathbf{r},t)-i\mu t}$ with the chemical potential $\mu$. The condensate flow velocities are given by $\mathbf{v}_{i}={\gradient}\theta_{i}/m_{a}$.
The equations for $\rho_{i}$ and $\theta_{i}$ of the condensate wave functions can be found in Visser and Weinfurtner (2005); Liberati et al. (2006); Syu et al. (2022). The perturbations around the stationary wave functions are defined through $\hat{\Psi}_{i}=\langle\hat{\Psi}_{i}\rangle(1+\hat{\phi}_{i})\,,$ where the fluctuation fields decompose in terms of the density and the phase as $\hat{\phi}_{i}=\delta{\hat{n}_{i}}+i\delta\hat{\theta}_{i}=\frac{\delta\hat{\rho}_{i}}{2\rho_{i}}+i\delta\hat{\theta}_{i}\,.$ According to Visser and Weinfurtner (2005), one can decouple the equations for the general spatially dependent condensate wave functions and the coupling strengths by choosing the time-independent background solutions $\rho_{1}=\rho_{2}=\rho$, $\theta_{1}=\theta_{2}=\theta$ or $\theta_{1}=\theta,\theta_{2}=\theta\pm\pi$, and $g_{11}=g_{22}=g$ (see Appendix. A) Zibold et al. (2010); Tommasini et al. (2003). The chosen scattering parameters of $g_{12}<g$ in the binary systems can result in a miscible state of the background condensates Kim et al. (2020); Hamner et al. (2011, 2013).
The decoupled equations are shown in Visser and Weinfurtner (2005); Liberati et al. (2006); Syu et al. (2022) in terms of the (gapless) density and (gapped) spin modes given by $\sqrt{2}\delta\hat{n}_{d/s}=\delta\hat{n}_{1}\pm\delta\hat{n}_{2}$ and $\sqrt{2}\delta\hat{\theta}_{d/s}=\delta\hat{\theta}_{1}\pm\delta\hat{\theta}_{2}$. Combining them yields a single equation for $\delta\hat{\theta}_{s}$ in the form of the Klein-Gordon equation for a massive scalar field,
@@
where the spatial dependent sound speed is given by $c^{2}=[(g-g_{12})\rho\pm\Omega]/m_{a}$, and the effective mass by $m^{2}_{\text{eff}}=\pm 2m_{a}^{2}c\,\Omega/\rho$, with the “$+$” sign corresponding to the zero-phase difference and the “--” sign to the $\pi$-phase difference. The acoustic metric $\mathbf{g}^{\mu\nu}$ depends on the background $\rho$ and $\theta$. Following Patrick et al. (2022), to mimic the geometry of a rotating black hole, we consider the draining vortex with the velocity flow
@@
where the winding number $\ell$ is taken as an integer (hereafter $\ell=1$) for quantum vortices and $d>0$ for the draining flow.
In the Thomas-Fermi approximations, the region of interest is far from the core of the vortex with the radius $r\gg R_{\text{TF}}=\sqrt{\frac{d^{2}+l^{2}}{2m_{a}\rho_{\infty}(g+g_{12})}}$, where the density $\rho$ can be treated as a constant Patrick et al. (2022), $\rho\simeq\rho_{\infty}$ given by $\rho_{\infty}=\frac{1}{g+g_{12}}\left(\mu-V_{0}\pm\frac{\Omega}{2}\right)$ for an uniform external potential $V_{0}$. The details can be seen in Appendix A. Using the coordinate transformations $d\tilde{t}=dt-(c^{2}/v_{r}^{2}-1)^{-1}dr,$ $d\tilde{\phi}=d\phi-\frac{v_{r}v_{\phi}}{r(c^{2}-v_{r}^{2})}dr,$ the metric becomes
@@
$\propto-\left(1-\frac{r_{E}^{2}}{r^{2}}\right)d\tilde{t}^{2}+\left(1-\frac{r^{2}_{H}}{r^{2}}\right)^{-1}dr^{2}$
$-\frac{2\ell}{m_{a}}\,d\tilde{\phi}d\tilde{t}+r^{2}d\tilde{\phi}^{2}\,$ (4)
with $\tilde{x}=(\tilde{t},r,\tilde{\phi})$. The acoustic horizon and the ergosphere are located at $r_{\text{H}}=d/m_{a}c$ and $r_{\text{E}}=\sqrt{d^{2}+\ell^{2}}/m_{a}c$, respectively. Hereafter, we simplify the formula by defining dimensionless variables as $r/r_{H}\rightarrow r$, $c\,\tilde{t}/r_{H}\rightarrow{t}$ and $r_{H}\Omega/c\rightarrow\Omega$. Recently, the collective gapped excitation mode was experimentally observed in the binary BECs Cominotti et al. (2022), prompting us to consider an experiment to simulate an analog black hole with environmental effects, which could be accessible by applying the space-dependent Rabi coupling Hamner et al. (2013); Mitsunaga et al. (2000); Nicklas et al. (2011); Han et al. (2015) in a two-dimensional BEC Fetter and Svidzinsky (2001). We consider the following parameters: the intraspecies scattering length $a_{11}=a_{22}=100a_{0}$, and the interspecies scattering length $a_{12}=90a_{0}$ , where $a_{0}$ is the Bohr radius. The atomic mass $m_{a}=1.44\times 10^{-25}\text{kg}$ and the asymptotic value of density $\rho_{\infty}$ can be prepared as much as $\sim 10^{21}\text{m}^{-3}$, giving the healing length and $R_{\text{TF}}$ such as $\xi\sim R_{\text{TF}}\sim 10^{2}\text{nm}$. Considering $r_{\text{H}}\sim 10\xi$, from $v_{r}=c$, the value of $d$ is about $d\gtrsim 1$. The corrections to the approximation of $\rho\simeq\rho_{\infty}$ are controlled by the small value of $R_{\text{TF}}/r_{H}$, which can be included beyond the constant density approximations in Oliveira et al. (2018); Demirkaya et al. (2020); Cardoso and Richartz (2022).
Also, the measuring time to trace the evolution of the perturbed fields will be limited by the lifetime of the condensates, say $\sim 1\text{s}$, giving the dimensionless time of order ${t}\sim 10^{3}$.
Using the phase fluctuation field of the form
$\delta\theta_{s}({t},r,\tilde{\phi})=\frac{H_{m}(r,{t})}{\sqrt{r}}\exp\left(im\tilde{\phi}\right)$ (5)
in the Klein-Gordon equation, the radial equation in the tortoise coordinate $r^{\ast}=r+\frac{1}{2}\log{\Big{|}\frac{r-1}{r+1}\Big{|}}$ ) reads
@@
×$\times\left[\frac{5}{4r^{4}}+\frac{1}{r^{2}}\left(m^{2}-\frac{1}{4}\right)+\mu_{s}^{2}(r)\right]\Bigg{\}}H_{m}(r,t)=0$ (6)
with $\mu^{2}_{s}(r)=\pm 2d\Omega(r)$. Further writing $H_{m}(r,t)=H_{\omega,m}(r)e^{-i\omega{t}}$ turns the time-dependent equation into the time-independent Schrodinger-like equation
$\left[\partial_{r^{\ast}}^{2}+V_{\text{eff}}(r)\right]H_{\omega,m}(r)=0,$ (7)
where
$V_{\text{eff}}(r)=$ $\left({\omega-\frac{m\ell}{r^{2}}}\right)^{2}-\left(\frac{r^{2}-1}{r^{2}}\right)$
×$\times\left[\frac{5}{4r^{4}}+\frac{1}{r^{2}}\left(m^{2}-\frac{1}{4}\right)+\mu^{2}_{s}(r)\right].$ (8)
Now the space-dependent mass term is introduced. The mass term due to the coupling to the surrounding matter of the black hole can be parametrized in terms of three parameters Cardoso et al. (2013a, b). One of them is the magnitude of the mass reflecting the coupling strength between the scalar field and matter in the scalar-tensor gravity theories, which can be controlled by the Rabi coupling $\Omega$. The other two are related to the parametrization of the mass distribution. They are the typical distance of the mass distribution $r_{0}$ from the black hole’s horizon and the distribution width $1/\alpha$. The parametrization of the distance $r_{0}$ can be attributed to the radius of the so-called innermost stable circular orbit of the particle moving around a massive object such as a black hole. The radius of the innermost circular motion depends on the mass and angular momentum (spin) of the massive object. The width of the mass distribution $1/\alpha$ is also an important parameter characterizing the surrounding matter of the black holes.
Compared to the counterpart of the effective potential defined from the Klein-Gordon equations in the scalar-tensor gravity theories, the existence of the ergosphere in the background metric is essential to account for the pure ingoing wave boundary condition at the horizon for $\omega<ml$. The presence of the mass term provides a potential barrier, leading to the possibility of quasibound states. Also, as $r\rightarrow\infty$, $V_{\text{eff}}(r)\rightarrow\omega^{2}$, where the pure outgoing wave solution can be achieved.
@@
Figure 2: In the case of positive mass squared, (a) shows the frequency for the mode $n=0$ (red) and $n=1$ (blue) as a function of $\Omega_{0}$ with fixed $r_{0}=15$, and (b) shows the corresponding frequency as a function of $r_{0}$ with $\Omega_{0}=0.2$. (c) depicts the effective potential and the normalized radial profile for the first two modes, and (d) presents a series of peaks at the resonant frequencies in the reflection coefficient spectrum. (d) reveals the growing waveform in time domain and the echo time duration. (c)–(e) share the same values of $\Omega_{0}$ and $r_{0}$ marked in the green line in (a) and (b). Note that $m=1$, and $\alpha=1/2$ are set in all figures.
Given the boundary conditions of the pure ingoing wave $e^{-i(\omega-m\ell)r^{\ast}}$ at the horizon and the pure outgoing wave $e^{i\omega r^{\ast}}$ at infinity, the eigenfrequency of the time-independent Schrodinger-like equation can be found with the complex values $\omega=\omega_{R}+i\omega_{I}$. However, there are two types of instabilities of interest with very different dynamics on the timescale $\tau\propto 1/|\omega_{I}|$: superradiant instabilities Patrick et al. (2018) are driven by an unstable but long-lived mode where $\omega_{R}\gg\omega_{I}$ , while tachyonic instabilities are featured with an almost pure imaginary frequency $\omega_{I}\gg\omega_{R}$ leading to a short-lived mode. In two-component BECs, the effective mass squared term can be parametrized as an analytically treatable form
@@
mimicking a mass shell surrounding the black hole Cardoso et al. (2013a, b); Cheung et al. (2022); Courty et al. (2023). The same instabilities are expected to be observed for a more general parametrization of the mass distribution, controlled by its typical distance from the black hole horizon and the width of the distribution.
However, there are other types of instabilities called ergoregion instabilities in Giacomelli and Carusotto (2020, 2021), which are not caused by a trap in the effective potential.
In the case of the spatially dependent mass term (9), we calculate the eigenfrequency numerically using the shooting method Cardoso and Yoshida (2005), where the solutions are obtained by integrating the equation from the horizon onward and from the far distance $r$ backward with appropriate boundary conditions where they continuously vary over $r$. We also find the eigenfrequency with the generalized Pöschl-Teller (PT) semianalytical methods Macedo et al. (2018); Torres et al. (2022); Casals et al. (2009). Additionally, the modified continued fraction method is adopted to obtain QNM’s frequencies shown in Appendix B Leaver (1990). The analytically approximated formulas for the eigenfrequency are given below. The instabilities are shown in the time domain by solving the time-dependent equation (6). For technical details, refer to Appendix C.
III Quasinormal modes and superradiant instabilities.
In the case of the quasinormal modes, their spectrum is strongly modified by the presence of the spatially dependent mass shell, which introduces an additional barrier to the rotational barrier given by the azimuthal angular momentum $m$. The frequencies of quasinormal modes for the real and imaginary parts are found in Fig. 1, which are smaller than their counterparts without the mass shell Berti et al. (2004); Cardoso et al. (2004a). The corresponding time domain signals are obtained by solving the time-dependent equation (6) for an initial Gaussian function with reference to Berti et al. (2022). According to Berti et al. (2022), small perturbations can lead to a relatively large shift of the eigenfrequency from the unperturbed one, but not to a significant change in the time domain (see an example in Appendix B). Here, we present the evolution of the radial profiles, which show a significant change due to the nonzero Rabi coupling strength. The echo time (estimated from the shaded time domain around $\sim 2r_{0}$), which is the time for the radial profile to travel towards the potential barrier at $r=r_{0}$ and return to the starting position and the ringdown of the profile at late times are displayed in Fig 1. More interestingly, when $\Omega_{0}$ or $r_{0}$ reaches its respective critical value, the positive mass shell could further induce superradiant instabilities, where a local bound state is shown for the potential ($-V_{\text{eff}}$) in Fig. 2. This is similar to superradiant instabilities induced by the accretion disk around Kerr black hole in scalar-tensor gravity theories Cardoso et al. (2013a, b); Lingetti et al. (2022). Let us estimate the resonance spectrum via a discrete quantization condition in the Wentzel-Kramers-Brillouin (WKB) analysis given by